OpenStudy (anonymous):

you have 500,000 dollars explain each scenerio and decide which is the best one to use?? Option 1: 6% compounded interest quarterly for 5 years. Option 2: 8% compounded interest annually for 5 years. Option 3: 14.5% simple interest for 10 years.

OpenStudy (anonymous):

For options 1 and 2, you'll use the formula A = P(1+r/n)^(n*t) For Option 1, you'll plug in P = 500000, r = 0.06, n = 4, and t = 5 For Option 2, you'll plug in P = 500000, r = 0.08, n = 1, and t = 5 ----------------------------------------------------------------- For option 3, you'll use the formula A = P(1+r*t) where in this case, P = 500000, r = 0.145, and t = 10

OpenStudy (anonymous):

so do I times all 3 options seperatly or what. In other words, how do I get the answer to all three options ? I am so lost, please help.